# Rope to Nope and Back Again: A New Hybrid Attention Strategy

### v1

@@ 3 Model Architecture
Based on the analysis above, we make the following design choices: the following design choices are made for the model architecture, building on top of the Command R+ architecture (Cohere For AI, 2024). First, we remove the QK-Norm component due to its poorly shaped attention patterns, which adversely impacted long context performance. Second, NoPE layers with a full attention span are introduced to enhance the model’s retrieval capabilities. Third, a sliding window size of 4,096 is applied to RoPE embeddings, leveraging RoPE’s inherent recency bias to improve performance on short-to-medium context ranges. In particular, the sliding-window approach has been employed in several prior works (Team et al., 2024a; Jiang et al., 2023; Character.AI, 2024). Regarding the number of layers, we perform an ablation study on the interleaving ratio of full attention and sliding window layers, testing the configurations of 1:1, 1:3, and 1:7. The results show that a 1:3 ratio strikes an optimal balance between computational efficiency and performance. Furthermore, we investigate the sequential ordering of full attention layers and find that it did not affect the results. As a result, we position the full attention layers at the end of each interleaving group, preceded by three sliding window layers. All other hyperparameters for the model architecture remain consistent with those outlined in Table 1.
Moreover, the combination of RoPE embeddings with sliding window layers effectively reduces computation cost and memory overhead while maintaining a recency bias, which is well suited for short-to-medium context ranges. Meanwhile, NoPE layers with full attention spans are employed to capture long-range dependencies, addressing the limitations of purely local architectures. Earlier transformer variants, such as GPT-J (Wang & Komatsuzaki, 2021) and GPT-NeoX (Black et al., 2022), also explored combining RoPE and NoPE by applying rotational embeddings to a portion of the head dimensions. However, the effectiveness of these approaches in length extrapolation remains an open question and an area of ongoing research.
For additional context, the resulting design shares similarities with other well-explored long-sequence architectures, such as Mega (Ma et al., 2023) and state-space models (SSMs) (Gu et al., 2022; Gu & Dao, 2024). For example, (Ma et al., 2023) and (Ma et al., 2024) introduced a multi-dimensional damped version of an exponential moving average component in conjunction with the gated attention unit (GAU) architecture (Hua et al., 2022), aiming to balance local and long-term dependencies, a common challenge in time-series modeling. Similarly, previous studies have proposed a diagonal variant of the S4 architecture (Gu et al., 2022), incorporating an exponentially decaying measure to enhance the model’s ability to capture long-range dependencies (Tay et al., 2020b).
@@ 4 Experiments
Pretraining and Cooldown. The models are pretrained for 5 trillion tokens of diverse data with batch size of 8 million tokens using FP8 precision format (Micikevicius et al., 2022). We use a cosine learning rate schedule of 5e-3 peak learning rate and 5% end learning rate with 8,000 linear warmup steps. From the pre-trained model, we linearly anneal the learning rate from 2.5e-4 to 1e-6 for 50,000 steps in BF16 precision. The context length was initially maintained at 8k for the first 35,000 steps, then extended to 32k and 128k for 10,000 steps and 5,000 steps, respectively. For the baseline model, the RoPE $\theta$ values were increased to 1,000,000 for 32k and 8,000,000 for 128k contexts during the length extrapolation phase, while remaining constant for the RNope-SWA variant. Both models utilized the same interleaved training strategy outlined in Section 2.1.
Supervised Finetuning. We supervise fine-tune on top of the pretrained models. As the primary focus of this study is to evaluate the impact of architectural design on downstream tasks, preference training is deferred to future work. To preserve the long context capabilities of the model, the fine-tuning process utilizes interleaved datasets containing 8k and 128k prompt-response pairs. The long context SFT data at 128k includes Retrieval-Augmented Generation (RAG) (Lewis et al., 2021) datasets, multilingual translation datasets with extended passages, long code instruction datasets, and long context retrieval datasets. Training was performed for two epochs across the entire dataset.
@@ 5 Experimental Results
Our evaluation contains a comprehensive analysis over standard benchmarks below 8k context length such as MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), ARC (Clark et al., 2018), SAT (Zhong et al., 2023), GSM8k (Cobbe et al., 2021), Winograde (Sakaguchi et al., 2019) and MBPP (Austin et al., 2021), as well as popular long context benchmarks with needles-in-a-haystack (Kamradt, 2023) and the retrieval and QA portion of Ruler (Hsieh et al., 2024). We test the context lengths up to 256k so we can examine the impact of these choices in the extrapolation capability of the model. We denote the baseline model trained with RoPE scaling as “Baseline” and the architecture with interleaved attention span and position embeddings as “RNope-SWA”.
@@ 5.1 Standard Benchmarks
Model MMLU HellaSwag ARC-E ARC-C SATEn SATMath GSM8K Winogrande MBPP
Baseline 57.5 75.8 84.6 48.5 70.0 30.9 40.9 68.5 39.1
RNope-SWA 59.5 76.2 82.5 48.8 71.9 30.5 42.7 69.5 39.3
@@
In this set of evaluations, we evaluate baseline and RNope-SWA on a standard LLM benchmark covering various language, math and code capabilities. The results are shown in Table 5. We observe that the model is better or on-par on most of the metrics compared to the baseline and has some gains over the baseline numbers on certain benchmarks (+2.0% on MMLU and +1.8% on GSM8k). This set of results also indicates that although RNope-SWA explicitly removed position embeddings from 25% of all its layers, positional information is retained by the interleaving attention span and captured by RoPE from previous layers. RNope-SWA does not have the performance loss due to the removal of explicit position embeddings, as previous works have shown (Kazemnejad et al., 2023; Wang et al., 2024).
@@ 5.2 Long Context Benchmarks
To evaluate the long context performance of these models, we use NIAH and the retrieval and QA components of Ruler (Hsieh et al., 2024). To better understand how architectural choices affect long context performance, we also evaluate with context lengths extending beyond training sequence length. This allows us to assess how well these models can interpolate as well as extrapolate to unseen context lengths and how specific architectural decisions influence these capabilities.
@@ 5.2.1 NIAH Evaluations
Following the settings of section 2.2, we run NIAH test twice with 128k and 256k context lengths respectively. The scores are reported in Figure 2. The figure indicates that although both models are able to get close to perfect scores below the context length seen during training, RNope-SWA has better extrapolation capabilities and achieves almost no loss up to 256k context length while the baseline fails to extrapolate well – despite using a high RoPE $\theta$ value of 8 million.
(a) Baseline model up to 128k. The raw score is 9.99
(b) RNope-SWA up to 128k. The raw score is 9.99
Figure 1: NIAH evaluation at 128k
(c) Baseline model up to 256k. The raw score is 8.25
(d) RNope-SWA up to 256k. The raw score is 9.97
Figure 2: NIAH evaluation at 128k and 256k
@@ 5.2.2 Ruler Evaluations
First introduced in (Hsieh et al., 2024), the Ruler benchmark aims to provide a set of more difficult tasks than NIAH. It covers a wider range of retrieval under a Multi-Query/Key/Value settings, more realistic tasks with a long-context Question-Answering format and more. Although our modification of NIAH introduced more context variants and proves to be more difficult than the vanilla version, it still cannot test the limits of the model. Therefore, we evaluate our models on the retrieval and QA portion of the Ruler so we can better separate their performance.
From the results, we can observe that the baseline model with RoPE $\theta$ scaling approach suffers from a sharp drop in the longer context range, especially 64k and longer. Comparing the difference between the scores obtained at 8k and 256k, the baseline model dropped from 96.6 to 57.1 (about 41%) on retrieval and from 53.5 to 30.0 (about 44%) on QA, whereas the RNope-SWA model dropped from 96.1 to 74.8 (about 22.1%) on retrieval and from 55.5 to 42.5 (about 23.4%) on QA respectively. From the original Ruler Paper (Hsieh et al., 2024), models that adopt similar RoPE scaling approaches have shown a similar degradation over longer context lengths (Dubey et al., 2024; Yang et al., 2024a; Abdin et al., 2024) as the baseline.
@@ 5.3 Impacts on Training and Inference
We also report the differences in training and inference speed, as well as memory requirements, of RNope-SWA compared to the baseline model. Let $S$ denote the sliding window size and $L$ represent the full training context length. During training, 75% of all layers now operate with a computational complexity of $\mathcal{O}(\textit{S}\textit{L})$, rather than the quadratic complexity of $\mathcal{O}(L^{2})$. This results in the model being approximately 50% faster than the baseline at a 64k context length and nearly 2x faster at 128k in terms of training throughput, using flash attention (Dao et al., 2022; Dao, 2023; Shah et al., 2024) and a sequence-parallel scheme similar to (Jacobs et al., 2023; Yang et al., 2024b). With alternative implementations, such as Ring Attention (Liu & Abbeel, 2023; Liu et al., 2023) or its variants (Brandon et al., 2023), sliding window adoption can reduce key-value block communication overhead if carefully sharded along the sequence dimension, potentially improving speed.
For inference, there is a theoretical upper bound of 75% for memory savings on kv cache and time complexity reductions. Empirically, we observed an approximate 44% reduction in end-to-end latency compared to the baseline model with full attention when using 132k input tokens and 96 output tokens. With 990k input tokens and 8 output tokens, the latency reduction increases to nearly 70%. As the sequence length increases, the latency savings approach the theoretical limit of 75% within our architectural setting. Furthermore, increasing the proportion of sliding window layers relative to full attention layers can further enhance both speed and memory efficiency.
@@ 6 Discussions and Future Work
Our findings align with recent work, such as YoCo (Sun et al., 2024), Jamba-1.5 (Team et al., 2024b), and MiniMax-01 (MiniMax et al., 2025), which demonstrate that hybrid attention mechanisms generally outperform full attention mechanisms in handling long contexts. However, the underlying reasons behind this seemingly counterintuitive observation remain largely unexplored. This opens an intriguing area of study, particularly as models push towards multi-million-token context lengths. Re-visiting and re-thinking the foundational components of transformer architectures, such as attention mechanisms, may become essential to accommodating these extreme requirements. Recent works (Ye et al., 2024; Leviathan et al., 2024; Soh et al., 2024) have begun to explore this direction by focusing on reducing attention noise across large context windows, a promising approach to refine the performance of attention modules.
In the mean time, there is growing interest in more sophisticated techniques for integrating contextual and positional information. These methods have significant potential but introduce challenges in maintaining runtime efficiency and compatibility with optimization frameworks like flash attention (Dao, 2023). Research efforts (Golovneva et al., 2024), have begun exploring these possibilities, offering new directions for advancing both efficiency and performance in long context processing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning algorithms. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Attention Distribution of all lengths
This table contains the attention distribution of RoPE and RNoPE variants over 8k, 32k and 128k sequence lengths.
Context Length Model NoPE Layers RoPE Layers
begin needle context qc begin needle context qc
8k RoPE - - - - 0.3863 0.0328 0.3809 0.2000
RNoPE-10k 0.3900 0.0952 0.4736 0.0412 0.1255 0.0102 0.5340 0.3302
RNoPE-100k 0.3854 0.0932 0.4783 0.0430 0.2135 0.0136 0.4558 0.3171
RNoPE-2M 0.3775 0.0902 0.4874 0.0449 0.2041 0.0126 0.4952 0.2881
RNoPE-4M 0.4153 0.0546 0.5072 0.0229 0.1389 0.0136 0.6162 0.2313
RNoPE-10k-swa 0.3830 0.1025 0.4702 0.0443 0.2040 0.0110 0.5938 0.1911
32k RoPE - - - - 0.3541 0.0201 0.4343 0.1915
RNoPE-10k 0.3275 0.0765 0.5672 0.0287 0.0049 0.0004 0.6805 0.3142
RNoPE-100k 0.3263 0.0778 0.5633 0.0327 0.0241 0.0005 0.6782 0.2972
RNoPE-2M 0.3250 0.0712 0.5735 0.0303 0.1111 0.0046 0.6233 0.2611
RNoPE-4M 0.3486 0.0369 0.5981 0.0165 .0960 0.0039 0.6774 0.2227
RNoPE-10k-swa 0.3303 .0742 0.5634 0.0321 - - - -
128k RoPE - - - - 0.3463 0.0010 0.4751 0.1776
RNoPE-10k 0.2991 0.0444 0.6430 0.0135 .0000 0.0001 0.7230 0.2769
RNoPE-100k 0.2454 0.0419 0.7016 0.0111 0.0001 0.0000 0.7749 0.2250
RNoPE-2M 0.2600 0.0401 0.6836 0.0162 .0417 0.0008 0.7516 0.2059
RNoPE-4M 0.2949 0.0307 0.6635 0.0109 0.0663 0.0022 0.7115 0.2230
RNoPE-10k-swa 0.2760 0.0467 0.6615 0.0159 - - - -
@@ Appendix B Attention Distribution of RoPE and QK-Norm variants
In this section, we present three plots comparing the attention distribution of RoPE and QK-Norm variants across sequence lengths of 8k, 32k, and 128k on needle samples, following the setup described in 2.2.2. Additionally, we provide the aggregated attention entropy for each variant to quantitatively validate the arguments made in 2.2.2.
To improve the clarity of the distribution plots, we preprocess the attention distribution array by removing the first 10 tokens and the last 3% of tokens at the end of each sequence. This step eliminates the disproportionate attention mass caused by the attention sink effect and the recency bias observed of RoPE, making the attention patterns more legible. We then compute a moving average with a window size of 100 tokens and average across all samples and layers to generate the final distributions.
@@
(a) Context Length 8k
(b) Context Length 32k
(c) Context Length 128k
Figure 3: Attention Distribution Across Sequence lengths

### v2

@@ 3 Model Architecture
Figure 3: RNope-SWA Model Architecture. SWA stands for sliding window attention.
Based on the analysis above, we make the following design choices on top of the Command R+ architecture [17]. First, we remove the QK-Norm component due to its poorly shaped attention patterns, which adversely impacted long context performance. Second, NoPE layers with a full attention span are introduced to enhance the model’s retrieval capabilities. Third, a sliding window size of 4,096 is applied to RoPE embeddings, leveraging RoPE’s inherent recency bias to improve performance on short-to-medium context ranges. In particular, the sliding-window approach has been employed in several prior works [67, 36, 11]. Regarding the number of layers, we perform an ablation study on the interleaving ratio of full attention and sliding window layers, testing the configurations of 1:1, 1:3, and 1:7. The results show that a 1:3 ratio strikes an optimal balance between computational efficiency and performance. We position the full attention layers at the end of each interleaving group, preceded by three sliding window layers. All other hyperparameters for the model architecture remain consistent with those outlined in Table 1. A visualization is shown in Figure 3.
For additional context, the resulting design shares similarities with other well-explored long-sequence architectures, such as Mega [48] and state-space models (SSMs) [28, 27]. For example, [48] and [49] introduced a multi-dimensional damped version of an exponential moving average component in conjunction with the gated attention unit (GAU) architecture [33], aiming to balance local and long-term dependencies, a common challenge in time-series modeling. Similarly, previous studies have proposed a diagonal variant of the S4 architecture [28], incorporating an exponentially decaying measure to enhance the model’s ability to capture long-range dependencies [65].
Earlier transformer variants, such as GPT-J [71] and GPT-NeoX [8], explored hybridizing RoPE and NoPE by applying rotational embeddings to a subset of the head dimensions. The work of p-RoPE [6] further advanced this line of research by analyzing RoPE in the frequency domain, revealing that low-frequency components capture semantic relationships while high-frequency components encode positional information. Similar to prior partial-RoPE approaches, p-RoPE removes selected low-frequency components along the head dimensions. These findings align with our observation that NoPE primarily facilitates retrieval (semantic focus), whereas RoPE exhibits a stronger recency bias (positional focus). However, prior works did not identify the cross-layer patterns underlying the “division of labor” phenomenon described in Section 2.2.3, which enables our approach to achieve both improved performance and practical engineering benefits.
@@ 4 Experiments
Pretraining and Cooldown. The models are pretrained for 5 trillion tokens of diverse data with batch size of 8 million tokens using FP8 precision format [51]. We use a cosine learning rate schedule of 5e-3 peak learning rate and 5% end learning rate with 8,000 linear warmup steps. From the pre-trained model, we linearly anneal the learning rate from 2.5e-4 to 1e-6 for 50,000 steps in BF16 precision. The context length was initially maintained at 8k for the first 35,000 steps, then extended to 32k and 128k for 10,000 steps and 5,000 steps, respectively. For the baseline model, the RoPE $\theta$ values were increased to 1,000,000 for 32k and 8,000,000 for 128k contexts during the length extrapolation phase, while remaining constant for the RNope-SWA variant. Both models utilized the same interleaved training strategy outlined in Section 2.1.
Supervised Finetuning. We supervise fine-tune on top of the pretrained models. As the primary focus of this study is to evaluate the impact of architectural design on downstream tasks, preference training is deferred to future work. To preserve the long context capabilities of the model, the fine-tuning process utilizes interleaved datasets containing 8k and 128k prompt-response pairs. The long context SFT data at 128k includes Retrieval-Augmented Generation (RAG) [40] datasets, multilingual translation datasets with extended passages, long code instruction datasets, and long context retrieval datasets. Training was performed for two epochs across the entire dataset.
@@ 5 Experimental Results
Our evaluation contains a comprehensive analysis over standard benchmarks below 8k context length such as MMLU [30], HellaSwag [79], ARC [15], SAT [82], GSM8k [16], Winograde [58] and MBPP [3], as well as popular long context benchmarks with needles-in-a-haystack [37] and the retrieval and QA portion of Ruler [32]. We test the context lengths up to 256k so we can examine the impact of these choices in the extrapolation capability of the model. We denote the baseline model trained with RoPE scaling as “Baseline" and the architecture with interleaved attention span and position embeddings as “RNope-SWA".
@@ 5.1 Standard Benchmarks
In this set of evaluations, we evaluate baseline and RNope-SWA on a standard LLM benchmark covering various language, math and code capabilities. The results are shown in Table 4. We observe that the model is better or on-par on most of the metrics compared to the baseline and has some gains over the baseline numbers on certain benchmarks (+2.0% on MMLU and +1.8% on GSM8k). This set of results also indicates that although RNope-SWA explicitly removed position embeddings from 25% of all its layers, positional information is retained by the interleaving attention span and captured by RoPE from previous layers. RNope-SWA does not have the performance loss due to the removal of explicit position embeddings, as previous works have shown [38, 72].
Model MMLU HellaSwag ARC-E ARC-C SATEn SATMath GSM8K Winogrande MBPP Baseline 57.5 75.8 84.6 48.5 70.0 30.9 40.9 68.5 39.1 RNope-SWA 59.5 76.2 82.5 48.8 71.9 30.5 42.7 69.5 39.3
@@ 5.2 Long Context Benchmarks
To evaluate the long context performance of these models, we use NIAH and the retrieval and QA components of Ruler [32]. To better understand how architectural choices affect long context performance, we also evaluate with context lengths extending beyond training sequence length. This allows us to assess how well these models can interpolate as well as extrapolate to unseen context lengths and how specific architectural decisions influence these capabilities.
@@ 5.2.1 NIAH Evaluations
Following the settings of section 2.2, we run NIAH test 256k context lengths. The scores are reported in Table 5 . The figure indicates that although both models are able to get close to perfect scores below the context length seen during training, RNope-SWA has better extrapolation capabilities and achieves almost no loss up to 256k context length while the baseline fails to extrapolate well – despite using a high RoPE $\theta$ value of 8 million. We attach the graphical visualization with depth and length dimension expanded in Appendix C.
Model Needles-128k Needles-256k Baseline 9.99 8.25 RNope-SWA 9.99 9.97
Table 5: Needles Score of Baseline and RNoPE-SWA up to 256k sequence length.
@@ 5.2.2 Ruler Evaluations
First introduced in [32], the Ruler benchmark aims to provide a set of more difficult tasks than NIAH. It covers a wider range of retrieval under a Multi-Query/Key/Value settings, more realistic tasks with a long-context Question-Answering format and more. Although our modification of NIAH introduced more context variants and proves to be more difficult than the vanilla version, it still cannot test the limits of the model. Therefore, we evaluate our models on the retrieval and QA portion of the Ruler so we can better separate their performance.
From the results, we can observe that the baseline model with RoPE $\theta$ scaling approach suffers from a sharp drop in the longer context range, especially 64k and longer. Comparing the difference between the scores obtained at 8k and 256k, the baseline model dropped from 96.6 to 57.1 (about 41%) on retrieval and from 53.5 to 30.0 (about 44%) on QA, whereas the RNope-SWA model dropped from 96.1 to 74.8 (about 22.1%) on retrieval and from 55.5 to 42.5 (about 23.4%) on QA respectively. From the original Ruler Paper [32], models that adopt similar RoPE scaling approaches have shown a similar degradation over longer context lengths [24, 75, 1] as the baseline.
@@ 5.3 Impacts on Training and Inference
We also report the differences in training and inference speed, as well as memory requirements, of RNope-SWA compared to the baseline model. Let $S$ denote the sliding window size and $L$ represent the full training context length. During training, 75% of all layers now operate with a computational complexity of $\mathcal{O}(\textit{S}\textit{L})$, rather than the quadratic complexity of $\mathcal{O}(L^{2})$. This results in the model being approximately 50% faster than the baseline at a 64k context length and nearly 2x faster at 128k in terms of training throughput, using flash attention [20, 19, 59] and a sequence-parallel scheme similar to [35, 76]. With alternative implementations, such as Ring Attention [43, 44] or its variants [9], sliding window adoption can reduce key-value block communication overhead if carefully sharded along the sequence dimension, potentially improving speed. In theory, KV cache memory and time complexity of RNoPE-SWA can be reduced by up to 75%. Empirically, we observed a 44% latency reduction at 132k input and 96 output tokens, and nearly 70% at 990k input and 8 output tokens—approaching the theoretical limit as sequence length grows. Increasing the ratio of sliding window to full attention layers can further improve speed and memory efficiency.
@@ 6 Discussions and Future Work
Our findings align with recent work, such as YoCo [62], Jamba-1.5 [68], and MiniMax-01 [52], which demonstrate that hybrid attention mechanisms generally outperform full attention mechanisms in handling long contexts. However, the underlying reasons behind this seemingly counterintuitive observation remain largely unexplored. This opens an intriguing area of study, particularly as models push towards multi-million-token context lengths. Re-visiting and re-thinking the foundational components of transformer architectures, such as attention mechanisms, may become essential to accommodating these extreme requirements. Recent works [77, 39, 60] have begun to explore this direction by focusing on reducing attention noise across large context windows, a promising approach to refine the performance of attention modules.
@@ Appendix A Attention Distribution of All Lengths
This table contains the attention distribution of RoPE and RNoPE variants from Section 2.2.2 over 8k, 32k and 128k sequence lengths.
Context Length Model NoPE Layers RoPE Layers Begin Needle Context End Begin Needle Context End 8k RoPE - - - - 0.3863 0.0328 0.3809 0.2000 RNoPE-10k 0.3900 0.0952 0.4736 0.0412 0.1255 0.0102 0.5340 0.3302 RNoPE-100k 0.3854 0.0932 0.4783 0.0430 0.2135 0.0136 0.4558 0.3171 RNoPE-2M 0.3775 0.0902 0.4874 0.0449 0.2041 0.0126 0.4952 0.2881 RNoPE-4M 0.4153 0.0546 0.5072 0.0229 0.1389 0.0136 0.6162 0.2313 RNoPE-10k-swa 0.3830 0.1025 0.4702 0.0443 0.2040 0.0110 0.5938 0.1911 32k RoPE - - - - 0.3541 0.0201 0.4343 0.1915 RNoPE-10k 0.3275 0.0765 0.5672 0.0287 0.0049 0.0004 0.6805 0.3142 RNoPE-100k 0.3263 0.0778 0.5633 0.0327 0.0241 0.0005 0.6782 0.2972 RNoPE-2M 0.3250 0.0712 0.5735 0.0303 0.1111 0.0046 0.6233 0.2611 RNoPE-4M 0.3486 0.0369 0.5981 0.0165 0.0960 0.0039 0.6774 0.2227 RNoPE-10k-swa 0.3303 0.0742 0.5634 0.0321 - - - - 128k RoPE - - - - 0.3463 0.0010 0.4751 0.1776 RNoPE-10k 0.2991 0.0444 0.6430 0.0135 0.0000 0.0001 0.7230 0.2769 RNoPE-100k 0.2454 0.0419 0.7016 0.0111 0.0001 0.0000 0.7749 0.2250 RNoPE-2M 0.2600 0.0401 0.6836 0.0162 0.0417 0.0008 0.7516 0.2059 RNoPE-4M 0.2949 0.0307 0.6635 0.0109 0.0663 0.0022 0.7115 0.2230 RNoPE-10k-swa 0.2760 0.0467 0.6615 0.0159 - - - -
@@ Appendix B Attention Distribution of RoPE and QK-Norm variants
In this section, we further investigate the suboptimal performance of the QK-Norm variant. We present three plots comparing the attention distribution between the RoPE and QK-Norm variants across sequence lengths of 8k, 32k, and 128k on needle samples, following the setup outlined in Section 2.2.2. Additionally, we provide the aggregated attention entropy for each variant to quantitatively support the arguments.
(a) Context Length 8k
(b) Context Length 32k
(c) Context Length 128k
Figure 4: Attention Distribution Across Sequence lengths
To enhance the clarity of the distribution plots, we preprocess the attention distribution array by removing the first 10 tokens and the last 3% of tokens from each sequence. This preprocessing step mitigates the disproportionate attention mass resulting from the attention sink effect and the recency bias observed in RoPE, thereby making the attention patterns more interpretable. We then compute a moving average with a window size of 100 tokens and average the results across all samples and layers to generate the final distributions.
@@
Appendix C Needles Score at 256k
(a) Baseline
(b) RNoPE-SWA
Figure 5: Needle Evaluation of Baseline and RNoPE-SWA on 256k sequence length